# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v1

@@ Abstract
Recent theoretical results show transformers cannot express sequential reasoning problems over long input lengths, intuitively because their computational depth is bounded. However, prior work treats the depth as a constant, leaving it unclear to what degree bounded depth may suffice for solving problems over short inputs, or how increasing the transformer’s depth affects its expressive power. We address these questions by analyzing the expressive power of transformers whose depth can grow minimally with context length $n$. We show even highly uniform transformers with depth $\Theta(\log n)$ can express two important problems: recognizing regular languages, which captures state tracking abilities, and graph connectivity, which underlies multi-step reasoning. Notably, both of these problems cannot be expressed by fixed-depth transformers under standard complexity conjectures, demonstrating the expressivity benefit of growing depth. Moreover, our theory quantitatively predicts how depth must grow with input length to express these problems, showing that depth scaling is more efficient than scaling width or chain-of-thought steps. Empirically, we find our theoretical depth requirements for regular language recognition match the practical depth requirements of transformers remarkably well. Thus, our results clarify precisely how depth affects transformers’ reasoning capabilities, providing potential practical insights for designing models that are better at sequential reasoning.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: To recognize a regular language over inputs of length $n$, the depth of a universal transformer can grow $\Theta(\log n)$ by Thm. 4.1. On the other hand, width must grow superpolynomially (Thm. 6.1), and the number of chain-of-thought steps must be superlogarithmic (Thm. 6.2). The precise depth and width coefficients plotted here were obtained experimentally in Section 7.
A line of recent work has analyzed the intrinsic computational power of transformers, the neural architecture behind today’s immensely successful large language models (LLMs). This work has established that, with fixed depth, transformers cannot express many simple problems outside the complexity class $\mathsf{TC}^{0}$, including recognizing regular languages and resolving connectivity of nodes in a graph (Merrill & Sabharwal, 2023a; Chiang et al., 2023). These problems conceivably underlie many natural forms of reasoning, such as state tracking (Liu et al., 2023; Merrill et al., 2024) or resolving logical inferences across long chains (Wei et al., 2022). Thus, these results suggest inherent limitations on the types of reasoning transformer classifiers can perform. Yet, while these results establish that transformers cannot solve these problems for arbitrarily long inputs, they come with an important caveat: transformers may still be able to solve such problems over inputs up to some bounded length, even if they cannot solve them exactly for inputs of arbitrary lengths. This perspective, coupled with the fact that treating depth as fixed is crucial to prior analyses placing transformers in $\mathsf{TC}^{0}$, motivates two related questions about depth as an important resource for a transformer, in relation to the context length over which it reasons:
@@
    Bounded Context: If fixed-depth transformers cannot theoretically express certain problems over unbounded context lengths, can they still express them over bounded but still practically “large enough” contexts? Can we quantitatively characterize transformers’ effective context length for different problems as a function of depth?
@@
    Dynamic Depth: Can minimally scaling the depth of a transformer allow it to solve such problems for arbitrarily long inputs? How does this compare in efficiency to scaling width or scaling inference-time compute via chain-of-thought steps (Merrill & Sabharwal, 2024)?
@@
We address these questions by analyzing the expressive power of “universal” transformers (also called “looped” transformers) where a fixed model is given dynamic depth by repeating a block of middle layers a variable number of times (Dehghani et al., 2019; Yang et al., 2024). We capture the regime where depth grows minimally with context length by allowing the middle layers to be repeated $\Theta(\log n)$ times on contexts of length $n$. We prove that such log-depth transformers can recognize regular languages and solve graph connectivity, two important reasoning problems known to be beyond fixed-depth transformers (Merrill & Sabharwal, 2023a). This has two interesting interpretations.
@@
First, this directly shows that, by dynamically increasing their depth as $\Theta(\log n)$ on inputs of length $n$, one can construct transformers that can solve regular language recognition (Thm. 4.1) and graph connectivity (Thm. 5.1) for arbitrary context length. In contrast, chain-of-thought steps, used for additional test-time compute by newest LLMs such as OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek AI, 2025), must be scaled superlogarithmically (Thm. 6.2) to solve these problems, and width must be scaled superpolynomially (Thm. 6.1), as shown in Fig. 1. Thus scaling depth more efficiently allows solving these reasoning problems compared to scaling width or using chain of thought.
@@
Second, a universal transformer unrolled to a fixed (independent of input size) depth $d$ is a special case of a standard $d$-depth transformer, namely one with a highly uniform structure (parameters are shared across layers). Thus, our result shows that standard transformers with a fixed depth $d$ can recognize regular languages (Cor. 4.3) and solve graph connectivity problems (Cor. 5.2) as long as one cares only about bounded inputs of size $2^{O(d)}$. This allows us to quantify how many layers are necessary for a desired input size. For instance, it follows from Corollaries 4.3 and 5.2 that with a depth of only $32$ (such as in smaller models like LLaMA 3.1 7B (Meta AI, 2024) and OLMo 7B (Ai2, 2024)), transformers can recognize regular languages up to strings of length $107$ and solve connectivity for graphs with up to $128$ vertices. A depth of $80$ (such as in LLaMA 3.1 70B) makes these input size limits practically unbounded: strings of length up to $440$K and graphs with up to $2.1$B vertices, respectively. Our experiments confirm that scaling depth as $\Theta(\log n)$ is necessary and sufficient for recognizing hard regular languages.
@@
Overall, we hope these findings serve as actionable guidance for practitioners to choose effective model depths for reasoning over long contexts, and motivate further exploration of dynamic depth as an inference-time compute strategy for transformer based LLMs.
@@ 2 Preliminaries: Universal Transformers
We consider $(s,r,t)$-universal transformers, which are defined to have $s$ fixed initial layers at the start, a sequence of $r$ layers that is repeated some number of times based on the input length, and a sequence of $t$ fixed final/terminal layers. Thus, an $(s,r,t)$-universal transformer unrolled $d(n)$ times for input length $n$ has a total of $s+rd(n)+t$ layers. A standard $d$-layer transformer is $(d,0,0)$-universal (equivalently, $(0,0,d)$-universal), while a standard universal transformer (Dehghani et al., 2019; Yang et al., 2024) is $(0,1,0)$-universal.
Definition 2.1.
@@
    An embedding projection matrix $\mathbf{E}:\Sigma\to\mathbb{Q}^{m}$ and positional encoding function $\pi:\mathbb{N}\to\mathbb{Q}^{m}$, which we assume separates $1$ from other indices (Merrill & Sabharwal, 2024); [^1] We use rationals $\mathbb{Q}$ instead of $\mathbb{R}$ so that the model has a finite description. All our simulations go through as long as at least $c\log n$ bits are used to represent rationals, similar in spirit to log-precision floats used in earlier analysis (Merrill & Sabharwal, 2023a, b).
@@
    A list of $s$ “initial” transformer layers (defined in Section 2.1);
@@
We next define how the transformer maps a sequence $w_{1}\cdots w_{n}\in\Sigma^{n}$ to an output value $y\in\Sigma$; to do so, we will always specify that the transformer is unrolled to a specific depth function $d(n)$, which we will consider to be $d(n)=\lceil\log n\rceil$. [^2] By computer science conventions, $\log n\triangleq\log_{2}n$. The computation is inductively defined by the residual stream $\mathbf{h}_{i}$: a cumulative sum of all layer outputs at each token $i$. In the base case, the residual stream $\mathbf{h}_{i}$ is initialized to $\mathbf{h}^{0}_{i}=\mathbf{E}(w_{i})+\pi(i)$. We then iteratively compute $s+rd(n)+t$ more layers, deciding which layer to use at each step as follows:
@@
We then compute $\mathbf{h}^{\ell}_{1},\ldots,\mathbf{h}^{\ell}_{n}=L^{\ell}(\mathbf{h}^{\ell-1}_{1},\ldots,\mathbf{h}^{\ell-1}_{n})$. The transformer output is a token determined by first computing the logits $\mathbf{h}^{\ell^{*}}_{n}\mathbf{U}$, where $\ell^{*}=s+rd(n)+t$, and then selecting the token with maximum score. We can identify special tokens in $\Sigma$ with “accept” and “reject” and define a transformer to recognize a language $L$ if, for every $w\in\Sigma^{*}$, it outputs “accept” if $w\in L$ and “reject” otherwise.
@@
2.1 Transformer Sublayers
To make Definition 2.1 well-defined, we will next describe the structure of the self-attention and feedforward sublayers that make up the structure of each transformer layer. Our definition of the transformer will have two minor differences from practice:
@@ 2.1 Transformer Sublayers
Each sublayer will take as input a sequence of normalized residual stream values $\mathbf{z}_{i}=\mathsf{layer\_norm}(\mathbf{m}\mathbf{h}_{i})$, where layer-norm can be standard layer-norm (Ba et al., 2016) or RMS norm (Zhang & Sennrich, 2019). The sublayer then maps $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$ to a sequence of updates to the residual stream $\delta_{1},\ldots,\delta_{n}$, and the residual stream is updated as $\mathbf{h}^{\prime}_{i}=\mathbf{h}_{i}+\delta_{i}$.
Definition 2.2 (Self-attention sublayer).
@@
$\mathbf{a}_{i,k}=\lim_{\alpha\to\infty}\sum_{j=1}^{c}\frac{\exp(\alpha\mathbf{q}_{i,k}^{\top}\mathbf{k}_{j,k})}{Z_{i,k}}\cdot\mathbf{v}_{j,k},$
$\;\textrm{where}\;Z_{i,k}=\sum_{j=1}^{c}\exp\left(\alpha\mathbf{q}_{i,k}^{\top}\mathbf{k}_{j,k}\right)$
and $c=i$ for causal attention and $c=n$ for unmasked attention. The $\alpha\to\infty$ limit implements averaging-hard attention: all probability mass is concentrated on the indices $j$ for which the attention score is maximized. This idealization is similar to assuming the temperature of the attention is large relative to the sequence length $n$. Finally, the attention heads are aggregated to create an output to the residual stream $\delta_{i}=\mathbf{W}\cdot\mathrm{concat}(\mathbf{a}_{i,1},\ldots,\mathbf{a}_{i,h})$.
Definition 2.3 (Feedforward sublayer).
@@ 2.2 Memory Management in Universal Transformers
Appendix A discusses in detail how $\{-1,0,1\}$ values can be stored directly in the residual stream, while a general scalar $z$ can be stored either as $\psi(z)=\langle z,1,-z,-1\rangle$ in its unnormalized form or as the unit vector $\phi(z)=\psi(z)/\sqrt{z^{2}+1}$ in its normalized form (cf. Merrill & Sabharwal, 2024). Importantly, however $z$ is stored, when it is read using masked pre-norm, we obtain $\phi(z)$. In Appendix A, we show how numerical values represented using $\psi$ or $\phi$ can be easily written (Lem. A.3), read (Lem. A.1), and deleted (Lemmas A.4 and A.5) from the residual stream. We will leverage these operations heavily in our theoretical constructions.
@@ 3 Fixed Depth Transformers Can Divide Small Integers
The overall idea is as follows. In the first layer, each position $i$ outputs an indicator of whether it’s a multiple of $m$. It also adds $\phi(j)$ to the residual stream such that $j$ is the quotient $i/m$ if $i$ is a multiple of $m$. In the second layer, each position $i$ attends to the nearest position $j\leq i$ that is a multiple of $m$ and retrieves the (normalized) quotient stored there, which is $j/m=\lfloor i/m\rfloor$. It adds this (normalized) quotient in its own residual stream. We then use Lem. A.6 to construct a third layer that adds $\phi(i-1)$ and $\phi(i-2)$ to the residual stream. A fourth layer checks in parallel whether the quotient stored at $i$ matches the quotients stored at $i-1$ and $i-2$, respectively. In the fifth layer, position $i$ counts the number of positions storing the same quotient as $i$, excluding the first such position. Finally, in the sixth layer, position $i$ attends to position $a_{i}$ to compute and add to the residual stream $\phi(\lfloor a_{i}/m\rfloor)$ (which is $\phi(b_{i})$) and $\phi(a_{i}-m\lfloor a_{i}/m\rfloor)$ (which is $\phi(c_{i})$). We next describe a detailed implementation of the construction, followed by an argument of its correctness.
@@
Construction. The first layer uses an attention head with queries, keys, and values computed as follows. The query at position $i$ is $q_{i}=\phi(i,m)=\phi(i/m)$ computed via Lem. A.2 leveraging the assumption that $\psi(i)$ and $\psi(m)$ are present in the residual stream. The key and value at position $j$ are $k_{j}=v_{j}=\phi(j)$. Let $h^{1}_{i}=\phi(j)$ denote the head’s output. The layer adds $h^{1}_{i}$ to the residual stream and also adds $e_{i}=\mathbb{I}(h^{1}_{i}=\phi(i/m))$ using Lem. A.7 (scalar equality check) on the first coordinate of $h^{1}_{i}$ and $\phi(i/m)$. As we will argue below, this layer has the intended behavior: $e_{i}=1$ if and only if $i$ is a multiple of $m$ and, if $e_{i}=1$, then the value it stores in the residual stream via $h^{1}_{i}$ is precisely the (normalized) quotient $i/m$ [^3] As described in Lem. A.7, a component will be added to the second layer to reset intermediate memory cells used in the first layer to 0 (this will happen analogously in later layers, but we will omit mentioning it).
@@
The second layer uses a head that attends with query $q_{i}=\langle 1,1\rangle$, key $k_{j}=\langle e_{j},[\phi(j)]_{0}\rangle$, and value $v_{j}=h^{1}_{j}$; note that both $e_{j}$ and $h^{1}_{j}$ can be read from the residual stream using masked pre-norm. This head attends to all positions $j\leq i$ that are multiples of $m$ (where $e_{j}=1$), with $[\phi(j)]_{0}$, the first component of $\phi(j)$, serving as a tie-breaking term for breaking ties in favor of the nearest multiple of $m$. Let $h^{2}_{i}=h^{1}_{j}$ denote the head’s output. The layer adds $h^{2}_{i}$ to the residual stream at position $i$. As we will argue below, $h^{2}_{i}=\phi(j/m)$ where $j/m$ is precisely the quotient stored in the residual stream at the multiple $j$ of $m$ that is closest to (and no larger than) $i$, which by definition is $\lfloor i/m\rfloor$. The layer thus adds $\phi(\lfloor i/m\rfloor)$ to the residual stream at position $i$.
The third layer uses Lem. A.6 to add $\phi(i-1)$ and $\phi(i-2)$ to the residual stream at $i$.
@@
In parallel for $k\in\{1,2\}$, the fourth layer attends with query $q_{i}=\phi(i-k)$, key $k_{j}=\phi(j)$, and value $v_{j}=\phi(\lfloor j/m\rfloor)$ to retrieve the quotient stored at position $i-k$. It uses Lem. A.7 (on the first coordinate) to store in the residual stream a boolean $b^{k}_{i}=\mathbb{I}(\phi(\lfloor i/m\rfloor)=\phi(\lfloor(i-k)/m\rfloor))$, indicating whether the quotient stored at $i$ matches the quotient stored at $i-k$.
In the fifth layer, position $i$ attends with query $q_{i}=\langle\phi(\lfloor i/m\rfloor),1\rangle$, key $k_{j}=\langle\phi(\lfloor j/m\rfloor),b^{1}_{j}\rangle$, and value $v_{j}=1-b^{2}_{j}$; note that $b^{k}_{i}$ can be retrieved from the residual stream. This head thus attends to every position with the same quotient as the current token besides the initial such position, with value 1 at the second such token and 0 elsewhere. Assuming $m$ does not divide $i$, this head will attend to precisely $i-m\lfloor i/m\rfloor$ positions and return $f_{i}=1/(i-m\lfloor i/m\rfloor)$ as the head output. The layer adds the vector $\psi(1,f_{i})$ defined as $\langle 1,f_{i},-1,-f_{i}\rangle$ to the residual stream at position $i$. This, when read in the next layer using masked pre-norm, will yield $\phi(1,f_{i})=\phi(1/f_{i})$. On the other hand, if $m$ does divide $i$ (which can be checked with a separate, parallel head), we write $\psi(0)$ to the residual stream, which, when read by the next layer, will yield $\phi(0)$.
The sixth layer attends with query $q_{i}=\phi(a_{i})$, key $k_{j}=\phi(j)$, and value $v_{j}=\langle h^{2}_{j},\phi(1/f_{j})\rangle$. Recall that $\phi(1/f_{j})$ can be read from the residual stream as discussed above. Further, the layer can recompute $f_{j}$ (or 0 in case $m$ divides $i$) and write $-\psi(1,f_{j})$ (or $-\psi(0)$, respectively) to the same coordinates, thereby resetting those cells to 0. Since $a_{i}\leq i$, the query matches exactly one position $j=a_{i}$, and the head retrieves $\langle h^{2}_{a_{i}},1/\phi(1/f_{a_{i}})\rangle$. This, by construction, is $\langle\phi(\lfloor a_{i}/m\rfloor),\phi(i-m\lfloor a_{i}/m\rfloor)\rangle$, which equals $\langle\phi(b_{i}),\phi(c_{i})\rangle$. The layer can thus store $\phi(b_{i})$ and $\phi(c_{i})$ to the residual stream at position $i$, as desired.
@@
The seventh and final layer cleans up any remaining intermediate values stored in the residual stream, setting them back to 0 as per Lem. A.7. This is possible because all values $v$ are of the form $\phi(x)$ or boolean, so adding $-\phi(v)$ will reset the cell to 0.
@@
We note that there are some high-level similarities between our division construction and a modular counting construction from Strobl et al. (2024), though the tools (and simplifying assumptions) used by each are different. Specifically, their approach relies on nonstandard position embeddings whereas ours makes heavy use of masked pre-norm.

### v2

@@ Abstract
Recent theoretical results show transformers cannot express sequential reasoning problems over long inputs, intuitively because their computational depth is bounded. However, prior work treats the depth as a constant, leaving it unclear to what degree bounded depth may suffice for solving problems over short inputs, or how increasing the transformer’s depth affects its expressive power. We address these questions by analyzing transformers whose depth can grow minimally with context length $n$. We show even highly uniform transformers with depth $\Theta(\log n)$ can express two important problems: recognizing regular languages, which captures state tracking abilities and was known to be expressible only by an unconventional, non-uniform model of transformers, and graph connectivity, which underlies multi-step reasoning. Notably, both of these problems cannot be expressed by fixed-depth transformers under standard complexity conjectures, demonstrating the expressivity benefit of growing depth. Moreover, our theory quantitatively predicts how depth must grow with input length to express these problems, showing that depth scaling is more efficient than scaling width or chain-of-thought steps. Empirically, our detailed experiments designed to bridge the expressivity vs. learnability gap reveal that our theoretical depth requirements for regular language recognition closely match the practical depth requirements for successfully training transformers. Thus, our results clarify how depth affects a transformer’s reasoning capabilities, and provide practical guidance for effective depth selection for sequential reasoning.
@@ 1 Introduction
A line of recent work analyzing the intrinsic computational power of transformers, the neural architecture behind today’s immensely successful large language models (LLMs), has established that, with fixed depth, transformers cannot express many simple problems outside the complexity class $\mathsf{TC}^{0}$, including recognizing regular languages and resolving connectivity of nodes in a graph (Merrill and Sabharwal, 2023a; Chiang et al., 2023). These problems conceivably underlie many natural forms of sequential reasoning, such as state tracking (Liu et al., 2023; Merrill et al., 2024) and resolving logical inferences across long chains (Wei et al., 2022). Thus, these results suggest inherent limitations on the types of reasoning transformer classifiers can perform. Yet, these findings come with an important caveat: even if transformers cannot solve such problems exactly for inputs of arbitrary lengths, they may still be able to solve them over inputs up to some bounded length. This perspective, coupled with the fact that treating depth as fixed is crucial to prior analyses placing transformers in $\mathsf{TC}^{0}$, motivates two related questions about depth as an important resource for a transformer, in relation to the context length over which it reasons:
@@
    Bounded Context: If fixed-depth transformers cannot theoretically express certain problems over unbounded context lengths, can they still express them over bounded but still practically “large enough” contexts? Can we quantitatively characterize the context length up to which transformers are effective for different problems as a function of their depth?
@@
    Dynamic Depth: Can minimally scaling the depth of a transformer allow it to solve such problems for arbitrarily long inputs? How does this compare in efficiency to scaling width (i.e., model dimension) or scaling inference-time compute via chain-of-thought steps?
@@
We address these questions by analyzing the expressive power of “universal” transformers (also called “looped” transformers) where a fixed model is given dynamic depth by repeating a block of middle layers a variable number of times (Dehghani et al., 2019; Yang et al., 2024; Geiping et al., 2025). We capture the regime where depth grows minimally with context length, with the middle layers repeated $\Theta(\log n)$ times on contexts of length $n$. We prove that even such highly uniform transformers, when allowed log-depth, can recognize regular languages [^1] While Liu et al. (2023) provided a similar result, their construction relied heavily on an unconventional, non-uniform transformer architecture, requiring a different set of model weights for each input length. In contrast, our result holds in a stronger setting of highly uniform transformers—where model weights are not only fixed (independent of input length) as in practice but even shared across blocks of layers, enabling effective inference-time scaling. Our formal model incorporates standard architectural choices like residual connection and layer norm. We supplement our stronger theoretical findings with matching empirical learnability results. and solve graph connectivity, two important reasoning problems known to be beyond fixed-depth transformers (Merrill and Sabharwal, 2023a). Our core technical contribution enabling this is Lemma 1, that fully uniform transformers can compute division and remainder of small integers. This not only obviates the need for non-uniformity and special positional encodings relied upon in prior work, it is also an interesting finding on its own.
Figure 1: To recognize a regular language over inputs of length $n$, the depth of a universal transformer can grow $\Theta(\log n)$ by Theorem 1. On the other hand, width must grow superpolynomially (Theorem 3), and the number of chain-of-thought steps must be superlogarithmic (Theorem 4). The precise depth and width coefficients plotted here were obtained experimentally in Section 7.
@@
Our result has two interesting interpretations. First, it directly shows that, by dynamically increasing their depth as $\Theta(\log n)$ on inputs of length $n$, one can construct transformers to solve regular language recognition (Theorem 1) and graph connectivity (Theorem 2) for arbitrary context length. In contrast, chain-of-thought (CoT) steps, used for additional test-time compute by newest LLMs such as OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek AI, 2025), must be scaled superlogarithmically (Theorem 4) to solve these problems, and width must be scaled superpolynomially (Theorem 3), as shown in Figure 1. Thus scaling depth more efficiently allows solving these reasoning problems compared to scaling width or using CoT.
@@
Second, a universal transformer unrolled to a fixed (independent of input size) depth $d$ is a special case of a standard $d$-depth transformer, namely one with a highly uniform structure (parameters are shared across layers). Thus, our result shows that standard transformers with a fixed depth $d$ can recognize regular languages (Corollary 1.2) and solve graph connectivity problems (Corollary 2.1) as long as one cares only about bounded inputs of size $2^{O(d)}$. This allows us to quantify how many layers are necessary for a desired input size. For instance, it follows from Corollaries 1.2 and 2.1 that with a depth of only $32$, such as in smaller models like LLaMA 3.1 7B (Meta AI, 2024) and OLMo 7B (Ai2, 2024), transformers can recognize regular languages up to strings of length $107$ and solve connectivity for graphs with up to $128$ vertices. A depth of $80$ (such as in LLaMA 3.1 70B) makes these input size limits practically unbounded: strings of length up to $440$K and graphs with up to $2.1$B vertices, respectively. Empirically, our experiments supplementing our theoretical findings demonstrate that scaling depth as $\Theta(\log n)$ is necessary and sufficient for learning to recognize hard regular languages.
@@
We hope these findings serve as actionable guidance for practitioners to choose effective model depths for reasoning over long contexts, and motivate further exploration of dynamic depth as an inference-time compute strategy for transformer based LLMs.
@@ 2 Preliminaries: Universal Transformers
We consider $(s,r,t)$-universal transformers, which are defined to have $s$ fixed initial layers at the start, a sequence of $r$ layers that is repeated some number of times based on the input length, and a sequence of $t$ fixed final/terminal layers. Thus, an $(s,r,t)$-universal transformer unrolled $d(n)$ times for input length $n$ has a total of $s+rd(n)+t$ layers. Geiping et al. (2025) empirically explored such transformers for scaling test-time computation for reasoning problems. A standard $d$-layer transformer is $(d,0,0)$-universal (equivalently, $(0,0,d)$-universal), while a standard universal transformer (Dehghani et al., 2019; Yang et al., 2024) is $(0,1,0)$-universal.
Definition 1.
@@
    An embedding projection matrix $\mathbf{E}:\Sigma\to\mathbb{Q}^{m}$ and positional encoding function $\pi:\mathbb{N}\to\mathbb{Q}^{m}$, which we assume separates $1$ from other indices (Merrill and Sabharwal, 2024); [^2] We use rationals $\mathbb{Q}$ instead of $\mathbb{R}$ so that the model has a finite description. All our simulations go through as long as at least $c\log n$ bits are used to represent rationals, similar in spirit to log-precision floats used in earlier analysis (Merrill and Sabharwal, 2023a, b).
@@
    A list of $s$ “initial” transformer layers (defined under “Transformer Sublayers” below);
@@
We next define how the transformer maps a sequence $w_{1}\cdots w_{n}\in\Sigma^{n}$ to an output value $y\in\Sigma$; to do so, we will always specify that the transformer is unrolled to a specific depth function $d(n)$, which we will consider to be $d(n)=\lceil\log n\rceil$. [^3] Following computer science conventions, $\log n\triangleq\log_{2}n$. The computation is inductively defined by the residual stream $\mathbf{h}_{i}$: a cumulative sum of all layer outputs at each token $i$. In the base case, the residual stream $\mathbf{h}_{i}$ is initialized to $\mathbf{h}^{0}_{i}=\mathbf{E}(w_{i})+\pi(i)$. We then iteratively compute $s+rd(n)+t$ more layers, deciding which layer to use at each step as follows:
@@
We then compute $\mathbf{h}^{\ell}_{1},\ldots,\mathbf{h}^{\ell}_{n}=L^{\ell}(\mathbf{h}^{\ell-1}_{1},\ldots,\mathbf{h}^{\ell-1}_{n})$. The transformer output is a token determined by first computing the logits $\mathbf{h}^{\ell^{*}}_{n}\mathbf{U}$, where $\ell^{*}=s+rd(n)+t$, and then selecting the token with maximum score. We can identify a special token in $\Sigma$ with “accept” and say that a transformer recognizes language $L$ if, for every $w\in\Sigma^{*}$, it outputs “accept” if and only if $w\in L$.
@@
Transformer Sublayers.
To make Definition 1 well-defined, we will next describe the structure of the self-attention and feedforward sublayers that make up the structure of each transformer layer. Our definition of the transformer will have two minor differences from practice:
@@
Each sublayer will take as input a sequence of normalized residual stream values $\mathbf{z}_{i}=\mathsf{layer\_norm}(\mathbf{m}\mathbf{h}_{i})$, where layer-norm can be standard layer-norm (Ba et al., 2016) or RMS norm (Zhang and Sennrich, 2019). The sublayer then maps $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$ to a sequence of updates to the residual stream $\delta_{1},\ldots,\delta_{n}$, and the residual stream is updated as $\mathbf{h}^{\prime}_{i}=\mathbf{h}_{i}+\delta_{i}$.
Definition 2 (Self-attention sublayer).
@@
$\mathbf{a}_{i,k}=\lim_{\tau\to 0}\sum_{j=1}^{c}\frac{\exp(1/\tau\cdot\mathbf{q}_{i,k}^{\top}\mathbf{k}_{j,k})}{Z_{i,k}}\cdot\mathbf{v}_{j,k},\quad\textrm{where}\;Z_{i,k}=\sum_{j=1}^{c}\exp\left(1/\tau\cdot\mathbf{q}_{i,k}^{\top}\mathbf{k}_{j,k}\right)$
and $c=i$ for causal attention and $c=n$ for unmasked attention. The $\tau\to 0$ limit implements averaging-hard attention: all probability mass is concentrated on the indices $j$ for which the attention score is maximized. This idealization is similar to assuming the temperature of the attention is large relative to the sequence length $n$. Finally, the attention heads are aggregated to create an output to the residual stream $\delta_{i}=\mathbf{W}\cdot\mathrm{concat}(\mathbf{a}_{i,1},\ldots,\mathbf{a}_{i,h})$.
Definition 3 (Feedforward sublayer).
@@ 2.1 Memory Management in Universal Transformers
Appendix A discusses in detail how $\{-1,0,1\}$ values can be stored directly in the residual stream, while a general scalar $z$ can be stored either as $\psi(z)=\langle z,1,-z,-1\rangle$ in its unnormalized form or as the unit vector $\phi(z)=\psi(z)/\sqrt{z^{2}+1}$ in its normalized form (“layer-norm hash”; cf. Merrill and Sabharwal, 2024). Importantly, however $z$ is stored, when it is read using masked pre-norm, we obtain $\phi(z)$. In Appendix A, we show how numerical values represented using $\psi$ or $\phi$ can be easily written (Lemma 4), read (Lemma 2), and deleted (Lemmas 5 and 6) from the residual stream. We will leverage these operations heavily in our theoretical constructions.
@@ 3 Fixed Depth Transformers Can Divide Small Integers
The overall idea is as follows. In the first layer, each position $i$ outputs an indicator of whether it’s a multiple of $m$. It also adds $\phi(j)$ to the residual stream such that $j$ is the quotient $i/m$ if $i$ is a multiple of $m$. In the second layer, each position $i$ attends to the nearest position $j\leq i$ that is a multiple of $m$ and retrieves the (normalized) quotient stored there, which is $j/m=\lfloor i/m\rfloor$. It adds this (normalized) quotient in its own residual stream. We then use Lemma 7A.3) to construct a third layer that adds $\phi(i-1)$ and $\phi(i-2)$ to the residual stream. A fourth layer checks in parallel whether the quotient stored at $i$ matches the quotients stored at $i-1$ and $i-2$, respectively. In the fifth layer, position $i$ counts the number of positions storing the same quotient as $i$, excluding the first such position. Finally, in the sixth layer, position $i$ attends to position $a_{i}$ to compute and add to the residual stream $\phi(\lfloor a_{i}/m\rfloor)$ (which is $\phi(b_{i})$) and $\phi(a_{i}-m\lfloor a_{i}/m\rfloor)$ (which is $\phi(c_{i})$). We next describe a detailed implementation of the construction, followed by an argument of its correctness.
@@
Construction. The first layer uses an attention head with queries, keys, and values computed as follows. The query at position $i$ is $q_{i}=\phi(i,m)=\phi(i/m)$ computed via Lemma 3A.1) leveraging the assumption that $\psi(i)$ and $\psi(m)$ are present in the residual stream. The key and value at position $j$ are $k_{j}=v_{j}=\phi(j)$. Let $h^{1}_{i}=\phi(j)$ denote the head’s output. The layer adds $h^{1}_{i}$ to the residual stream and also adds $e_{i}=\mathbb{I}(h^{1}_{i}=\phi(i/m))$ using Lemma 8 (scalar equality check, §A.4) on the first coordinate of $h^{1}_{i}$ and $\phi(i/m)$. As we will argue in Appendix B, this layer has the intended behavior: $e_{i}=1$ if and only if $i$ is a multiple of $m$ and, if $e_{i}=1$, then the value it stores in the residual stream via $h^{1}_{i}$ is precisely the (normalized) quotient $\phi(i/m)$. [^4] As described in Lemma 8, a component will be added to the second layer to reset intermediate memory cells used in the first layer to 0 (this will happen analogously in later layers, but we will omit mentioning it).
@@
The second layer uses a head that attends with query $q_{i}=\langle 1,1\rangle$, key $k_{j}=\langle e_{j},[\phi(j)]_{0}\rangle$, and value $v_{j}=h^{1}_{j}$; note that both $e_{j}$ and $h^{1}_{j}$ can be read from the residual stream using masked pre-norm. This head attends to all positions $j\leq i$ that are multiples of $m$ (where $e_{j}=1$), with $[\phi(j)]_{0}$, the first component of $\phi(j)$, serving as a tie-breaking term for breaking ties in favor of the nearest multiple of $m$. As we will argue in Appendix B, this head outputs $\phi(\lfloor i/m\rfloor)$, which we store in the residual stream.
The third layer uses Lemma 7A.3) to add $\phi(i-1)$ and $\phi(i-2)$ to the residual stream at position $i$.
@@
In parallel for $k\in\{1,2\}$, the fourth layer attends with query $q_{i}=\phi(i-k)$, key $k_{j}=\phi(j)$, and value $v_{j}=\phi(\lfloor j/m\rfloor)$ to retrieve the quotient stored at position $i-k$. It uses Lemma 8A.4) on the first coordinate to store in the residual stream a boolean $b^{k}_{i}=\mathbb{I}(\phi(\lfloor i/m\rfloor)=\phi(\lfloor(i-k)/m\rfloor))$, indicating whether the quotient stored at $i$ matches the quotient stored at $i-k$.
In the fifth layer, position $i$ attends with query $q_{i}=\langle\phi(\lfloor i/m\rfloor),1\rangle$, key $k_{j}=\langle\phi(\lfloor j/m\rfloor),b^{1}_{j}\rangle$, and value $v_{j}=1-b^{2}_{j}$. When the output of this layer is passed through layer norm, it produces $\phi(i\bmod m)$ as justified in Appendix B.
The sixth layer attends with query $q_{i}=\phi(a_{i})$, key $k_{j}=\phi(j)$, and value $v_{j}=\langle\lfloor j/m\rfloor,\phi(j\bmod m)\rangle$ (from layers two and five) to compute $\langle\phi(\lfloor a_{i}/m\rfloor),\phi(a_{i}\bmod m)\rangle$, which equals $\langle\phi(b_{i}),\phi(c_{i})\rangle$.
@@
The seventh and final layer cleans up any remaining intermediate values stored in the residual stream, setting them back to 0 as per Lemma 8. This is possible because all values $v$ are of the form $\phi(x)$ or a boolean, which means adding $-\phi(v)$ to the residual stream will reset the corresponding cell to 0.
@@
Our division construction is somewhat similar to the modular counting construction from Strobl et al. (2024), though the tools and underlying assumptions are different. Specifically, their approach relies on nonstandard position embeddings whereas ours uses masked pre-norm.